Explain why ESR frequencies are of the
order of GHz but NMR frequencies are of the order of MHz for the same applied magnetic field?
1
Expert's answer
2019-06-06T05:18:23-0400

Solution.

"h \times \omega = gl \times \mu \times H"

In the case of NMR, instead of the Bohr magneton and the gl - Lande factor, there is a nuclear magneton (NM) and a gc - is nuclear factor taking into account the complex structure of the nucleus. Due to the fact that M = 1836 * m, the resonance frequency of NMR is noticeably lower than the EPR frequency. Therefore, electron resonance is observed at microwave frequencies in the range of 1-10 Hz, whereas in the study of NMR, the range of short radio waves is usually used. Thus, NMR takes into account the structure of the nucleus, and the EPR motion of the electron.

Answer:

"h \times \omega = gl \times \mu \times H"

In the case of NMR, instead of the Bohr magneton and the gl - Lande factor, there is a nuclear magneton (NM) and a gc - is nuclear factor taking into account the complex structure of the nucleus. Due to the fact that M = 1836 * m, the resonance frequency of NMR is noticeably lower than the EPR frequency. Therefore, electron resonance is observed at microwave frequencies in the range of 1-10 Hz, whereas in the study of NMR, the range of short radio waves is usually used. Thus, NMR takes into account the structure of the nucleus, and the EPR motion of the electron.